# Single-loop Projection-free and Projected Gradient-based Algorithms for Nonconvex-concave Saddle Point Problems with Bilevel Structure

### v1

@@ 5.1 Convergence Rate of Algorithm 1
Assuming $L^{\Phi}_{yy}=0$, then, $\rho_{D}=0$ and $\sigma=\mathcal{O}(\frac{1}{\mu})$. By revisiting Lemmas 4.5, 4.6, and 4.7 we obtain $\Gamma_{1}=\mathcal{O}(\frac{\kappa_{g}}{\mu}L_{y\theta}^{\Phi})$, $\Gamma_{2}=\mathcal{O}\left(\frac{\kappa_{g}L_{y\theta}^{\Phi}+L_{yx}^{\Phi}}{\mu}\gamma+\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}}{\mu}\gamma\right)$, and $\Gamma_{3}=\mathcal{O}(\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}+\kappa_{g}L_{yx}^{\Phi}}{\mu}\gamma+\frac{\kappa_{g}^{3}L_{y\theta}^{\Phi}}{\mu}\gamma)$. Similar to the proof of Corollary 5.2, by updating $B_{2}=\mathcal{O}(\frac{\kappa_{g}^{4}}{\mu})$ and selecting $\mu=\mathcal{O}(1/(K)^{1/3})$, $\gamma=\mathcal{O}(1/(K)^{2/3})$, Algorithm 1 requires $K=\mathcal{O}(\frac{\kappa_{g}^{4}}{\epsilon^{3}})$ iterations to achieve $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\epsilon$.
@@
Our results in Theorem 5.1 represent a novel bound for the constrained nonconvex-concave saddle point problem, subject to a strongly convex problem in the lower-level setting. To the best of our knowledge, there has been no exploration of any projection-free algorithm alternative in this setup.
@@ 5.2 Convergence Rate of Algorithm 2
Suppose Assumptions 2.1, 2.2 hold and let $\{(x_{k},y_{k},\theta_{k})\}_{k\geq 0}$ be the sequence generated by Algorithm 2 with step-sizes $\gamma_{k}=\gamma>0$, $\eta_{k}=\eta>0$, $\alpha>0$, $\tau_{k}=\tau>0$, and $\sigma_{k}=\sigma\leq\frac{2}{L^{\Phi}_{yy}+2\mu}$ and regularization parameter $\mu>0$. Then, for any $K\geq\bar{K}=\mathcal{O}(\mu\log(\kappa_{g}/\mu))$ we have that
$\frac{1}{\tau^{2}K}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}$ $\leq\frac{2(f(x_{0})-f(x^{*}))+\mu D_{\mathcal{Y}}}{\tau\gamma K}+\frac{\gamma B_{4}}{\tau K},$ (63)
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|}{\sigma(1-\rho_{D})K}+\gamma B_{3}D_{\mathcal{X}}+\frac{3}{(1-\rho_{D})(1-\beta)}\frac{L_{y\theta}^{\Phi}}{K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mu D_{\mathcal{Y}},$ (64)
for some constant $B_{3},B_{4}>0$ where $B_{3}$ is defined in Theorem 5.1 and $B_{4}=\mathcal{O}\big{(}(\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}+\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}+\kappa_{g}^{4}\big{)}$.
@@
$f_{\mu}(x_{k+1})$ $\leq f_{\mu}(x_{k})+\langle\nabla f_{\mu}(x_{k})-\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k+1}-x_{k}\rangle+\langle\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x},x_{k+1}-x_{k}\rangle$
@@
$\quad+\langle G_{k}^{x},x_{k+1}-x_{k}\rangle+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\|{x_{k+1}-x_{k}}\|^{2}.$ (65)
@@
Recall the update of $s_{k}=\mathcal{P}_{\mathcal{X}}(x_{k}-\tau G_{k}^{x})$. The optimality condition of this step implies that
@@
$\left\langle G_{k}^{x},x_{k+1}-x_{k}\right\rangle=\gamma\left\langle G_{k}^{x},s_{k}-x_{k}\right\rangle\leq-\frac{\gamma}{\tau}\left\|s_{k}-x_{k}\right\|^{2}.$ (66)
Combining (66) with (5.2) and using $\|x_{k+1}-x_{k}\|=\gamma\left\|s_{k}-x_{k}\right\|$ we obtain
@@
$\frac{\gamma}{2\tau}\left\|s_{k}-x_{k}\right\|^{2}$ $\leq f_{\mu}(x_{k})-f_{\mu}(x_{k+1})+\gamma\mathbf{L}_{\mathcal{L}_{2}}\left\|y_{k}-y_{\mu}^{*}(x_{k})\right\|\left\|s_{k}-x_{k}\right\|+\gamma\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x}\|\left\|s_{k}-x_{k}\right\|$
$\quad+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma^{2}\left\|s_{k}-x_{k}\right\|^{2}-\frac{\gamma}{2\tau}\left\|s_{k}-x_{k}\right\|^{2}.$ (67)
Summing the above inequality over $k=0$ to $K-1$, dividing both sides by $\gamma K$, and using Lemmas 4.6, 4.8 we obtain
@@
$\frac{1}{2\tau K}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}$ $\leq\frac{f_{\mu}(x_{0})-f_{\mu}(x_{K})}{\gamma K}+\frac{\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+2\mathbf{C}_{v2}\Gamma_{1}+\Gamma_{4}}{K}$
$\quad+\left(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{2}+\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma-\frac{1}{2\tau}\right)\frac{1}{K}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}.$ (68)
@@
Selecting step-size $\tau$ such that $\tau\leq 2/(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{2}+\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma)$ we can drop the last term in RHS of the above inequality. Therefore, using the fact that $f_{\mu}(x_{0})-f_{\mu}(x_{K})\leq f(x_{0})-f(x^{*})+\frac{\mu}{2}D_{\mathcal{Y}}^{2}$ and multiplying both sides by $2/\tau$ the desired result can be obtained.
@@
Next, using the same line of proof as in Theorem 5.1 we can derive a bound for the dual gap function as in (5.1). Therefore, using boundedness of constraint set $\mathcal{X}$ and by noting that for any $k\geq 2$ $\sum_{i=1}^{k}\rho_{D}^{k-i}\beta^{i}\leq(\frac{\rho_{D}^{k/2}}{1-\beta}+\frac{\beta^{k/2}}{1-\rho_{D}})\beta$, we conclude that for any $k\geq 2$,
@@
Under the premises of Theorem 5.4, let $\mu=\mathcal{O}(\epsilon)$, $\gamma=\mathcal{O}(\epsilon^{3})$ and $\tau\leq\mathcal{O}(\epsilon(\kappa_{g}^{2}L_{y\theta}^{\Phi}(\kappa_{g}L_{y\theta}^{\Phi}+L_{yx}^{\Phi}))^{-1})$, then for any $K\geq\bar{K}=\mathcal{O}(\mu\log(\kappa_{g}/\mu))$, there exists $t\in\{\bar{K},\ldots,K-1\}$ such that $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$ satisfies $\mathcal{G}_{\mathcal{Z}}(x_{t},y_{t})\leq\epsilon$ within $\mathcal{O}(\frac{\kappa_{g}^{3}}{\epsilon^{5}}\log(\frac{\kappa_{g}}{\epsilon})+\frac{\kappa_{g}^{8}}{\epsilon^{2.5}}+\frac{\kappa_{g}^{6}}{\epsilon^{2}})$ iterations.
@@
Considering the result in (5.2) of Theorem 5.4 and multiplying both sides by $K/(K-\bar{K})$ for some $0<\bar{K}<K$. Then, we have that for any $K\geq 2$,
$\frac{1}{\tau^{2}(K-\bar{K})}\sum_{k=\bar{K}}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}$ $\leq\frac{1}{\tau^{2}(K-\bar{K})}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}$
$\leq\frac{2(f(x_{0})-f(x^{*}))+\mu D_{\mathcal{Y}}^{2}}{\tau\gamma(K-\bar{K})}+\frac{2(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+2\mathbf{C}_{v2}\Gamma_{1}+\Gamma_{4})}{\tau(K-\bar{K})}$
Let $k^{*}=\operatorname*{\arg\!\min}_{k=\bar{K}}^{K-1}\left\|s_{k}-x_{k}\right\|$. Lower-bounding the LHS and taking the square root from both sides of the above inequality, we obtain
$\mathcal{G}_{\mathcal{X}}(x_{k^{*}},y_{k^{*}})=\frac{1}{\tau}\left\|s_{k^{*}}-x_{k^{*}}\right\|$ $\leq\sqrt{\frac{2(f(x_{0})-f(x^{*}))+\mu D_{\mathcal{Y}}^{2}}{\gamma\tau(K-\bar{K})}}+\sqrt{\frac{2(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+2\mathbf{C}_{v2}\Gamma_{1}+\Gamma_{4})}{\tau(K-\bar{K})}}$
$=\mathcal{O}\left(\frac{1}{\sqrt{\gamma\tau K}}+\sqrt{\frac{\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu}{\tau K}}\right),$ (70)
where the last equality follows from $\Gamma_{1}=\mathcal{O}(\frac{\kappa_{g}}{\mu}L_{y\theta}^{\Phi})$, $\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$, and $\mathbf{L}_{\mathcal{L}_{2}}=\mathcal{O}(\kappa_{g})$.

### v3

@@ 5.1 Convergence Rate of Algorithm 1
Recall that assuming $L^{\Phi}_{yy}=0$, we have $\rho_{D}=0$, leading to simplification of the bounds in Lemmas 4.5, 4.6, and 4.7. In particular, we derive $\Gamma_{1}=\frac{L^{\Phi}_{y\theta}}{\mu(1-\beta)}\|\theta_{0}-\theta^{*}(x_{0})\|D_{\mathcal{X}}=\mathcal{O}(\frac{\kappa_{g}L^{\Phi}_{y\theta}}{\mu})$, $\Gamma_{2}=\frac{L^{\Phi}_{y\theta}\gamma\mathbf{L}_{\theta}\beta}{\mu(1-\beta)}=\mathcal{O}(\frac{\kappa_{g}^{2}L^{\Phi}_{y\theta}}{\mu}\gamma)$ and $\Gamma_{3}=\frac{2\Gamma_{2}}{1-\rho}+\frac{L_{y\mu}\gamma}{(1-\rho)\mu}=\mathcal{O}(\frac{\kappa_{g}^{3}L^{\Phi}_{y\theta}}{\mu}\gamma+\frac{\kappa_{g}^{2}L^{\Phi}_{y\theta}+\kappa_{g}L^{\Phi}_{yx}}{\mu}\gamma)$, simplifying the bound on the primal and dual gap functions obtained in Theorem 5.1 with the following constants $B_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$, $B_{2}=\mathcal{O}(\frac{\kappa_{g}^{4}}{\mu}(L^{\Phi}_{y\theta})^{2}+\kappa_{g}^{4})$, and $B_{3}=\mathcal{O}(\kappa_{g}^{2}L^{\Phi}_{y\theta}\gamma)$. Using a similar argument as in the proof of Corollary 5.2, and noting that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\mathcal{O}(\frac{1}{\gamma K}+\gamma B_{2}+\mu)$, choosing $\mu=\kappa_{g}^{2}/K^{1/3}$ and $\gamma=1/(\kappa_{g}K)^{2/3}$, it follows that $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\mathcal{O}(\frac{\kappa_{g}^{4/3}}{K^{1/3}})$. Therefore, to achieve $\mathcal{G}_{\mathcal{Z}}(x_{k^{*}},y_{k^{*}})\leq\epsilon$, Algorithm 1 requires $K=\mathcal{O}(\frac{\kappa_{g}^{4}}{\epsilon^{3}})$ iterations.
@@
Our results in Theorem 5.1 represent a new bound for the constrained nonconvex-concave saddle point problem subject to a strongly convex problem in the lower-level setting. Specifically, it introduces an upper bound on a gap function for the primal part of Problem (1), which corresponds to the linear minimization oracle (LMO), and an upper bound on a gap function for the dual part of Problem (1) for the iterates generated by Algorithm 1. To the best of our knowledge, this is the only one-sided projection-free algorithm for solving the general non-bilinear SPB problem. Additionally, this method reduces the computational cost compared to projection-based algorithms for certain constraint sets $\mathcal{X}$, such as the nuclear-norm ball, Schatten-norm ball, and certain polyhedron sets [25].
Remark 5.2.
One can replace the projection operator in the $y$-update with a proximal map of a proper lower semi-continuous convex function $h$ with a bounded domain and Algorithm 1 and convergence results in Corollaries 5.2 and 5.3 will still hold. Moreover, it is worth noting that in the case $L^{\Phi}_{yy}=0$, we can solve a bilevel optimization problem with a composite objective function of the form: $\min_{x}h(\tilde{\Phi}(x,\theta(x))$ s.t. $\theta(x)\in\operatorname*{\arg\!\min}_{\theta}g(x,\theta)$ where $\tilde{\Phi}$ is a smooth map. In this case, $\Phi(x,y)=\langle\tilde{\Phi}(x,\theta(x),y\rangle$, and based on Corollary 5.3, the complexity of achieving $\epsilon$-stationary solution is $\mathcal{O}(\frac{\kappa_{g}^{4}}{\epsilon^{3}})$, which constitutes a new result in the context of bilevel optimization.
@@ 5.2 Convergence Rate of Algorithm 2
In this subsection, we establish upper bounds for the primal and dual gap functions for the iterates generated by Algorithm 2. Additionally, we explore the iteration complexity of the i-BRPD:FP method when the upper-level objective function exhibits linearity in $y$ (i.e., $L^{\Phi}_{yy}=0$).
@@
Suppose Assumptions 2.1, 2.2 hold and let $\{(x_{k},y_{k},\theta_{k})\}_{k\geq 0}$ be the sequence generated by Algorithm 2 with step-sizes $\gamma_{k}=\gamma>0$, $\eta_{k}=\eta>0$, $\alpha>0$, $\tau_{k}=\tau>0$, and $\sigma_{k}=\sigma\leq\frac{2}{L^{\Phi}_{yy}+2\mu}$ and regularization parameter $\mu>0$. Then, for any $K\geq 1$, we have that
$\frac{1}{\tau^{2}K}\sum_{k=0}^{K-1}\left\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{s}_{k}}-x_{k}\right\|^{2}$ $\leq\frac{2(f(x_{0})-f(x^{*}))+\mu D_{\mathcal{Y}}}{\tau\gamma K}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+4\tau M\Gamma_{4}}{\tau K}+4\tau\gamma M^{2}B_{4}},$
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|}{\sigma(1-\rho_{D})K}+\gamma B_{3}D_{\mathcal{X}}+\Big{(}1+\frac{2}{1-\rho_{D}}\Big{)}\frac{L_{y\theta}^{\Phi}}{(1-\beta)K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mu D_{\mathcal{Y}},$
for some constant $B_{3},B_{4}>0$, where $B_{3}$ is defined in Theorem 5.1 and $B_{4}=\mathcal{O}(\kappa_{g}^{3}/\mu)$.
@@
$f_{\mu}(x_{k+1})$ $\leq f_{\mu}(x_{k})+\langle\nabla f_{\mu}(x_{k})-\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k+1}-x_{k}\rangle+\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k+1}-x_{k}\rangle$
$\quad+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\|{x_{k+1}-x_{k}}\|^{2}$
$=f_{\mu}(x_{k})+\langle\nabla f_{\mu}(x_{k})-\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k+1}-x_{k}\rangle+\gamma\langle G_{k}^{x},s_{k}-x_{k}\rangle$
@@
$\quad+\gamma\langle\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x},s_{k}-x_{k}\rangle+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\|{x_{k+1}-x_{k}}\|^{2},$ (26)
@@
where in the last inequality we add and subtract $\gamma\left\langle G_{k},s_{k}-x_{k}\right\rangle$. Recall the update of $s_{k}=\mathcal{P}_{\mathcal{X}}(x_{k}-\tau G_{k}^{x})$ and $x_{k+1}=\gamma s_{k}+(1-\gamma)x_{k}$. The optimality condition of this step implies that
$\gamma\left\langle G_{k}^{x},s_{k}-x_{k}\right\rangle=\left\langle G_{k}^{x},x_{k+1}-x_{k}\right\rangle\leq-\frac{\gamma}{\tau}\left\|s_{k}-x_{k}\right\|^{2}.$ (27)
Moreover, let us define $\tilde{s}_{k}=\mathcal{P}_{\mathcal{X}}(x_{k}-\tau\nabla_{x}\mathcal{L}(x_{k},y_{k}))$ which implies that $\mathcal{G}_{\mathcal{X}}(x_{k},y_{k})=\frac{1}{\tau}\|\tilde{s}_{k}-x_{k}\|$. Considering $\gamma\left\langle G_{k}^{x},s_{k}-x_{k}\right\rangle$ once again and adding and subtracting $\left\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),\tilde{s}_{k}-s_{k}\right\rangle$ we obtain
@@
$\gamma\left\langle G_{k}^{x},s_{k}-x_{k}\right\rangle$ $=\gamma\left\langle G_{k}^{x},s_{k}-\tilde{s}_{k}\right\rangle+\gamma\left\langle G_{k}^{x}-\nabla_{x}\mathcal{L}(x_{k},y_{k}),\tilde{s}_{k}-x_{k}\right\rangle+\gamma\left\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),\tilde{s}_{k}-x_{k}\right\rangle$
$\leq\gamma\left(\|G_{k}^{x}\|\|s_{k}-\tilde{s}_{k}\|+\|G_{k}^{x}-\nabla_{x}\mathcal{L}(x_{k},y_{k})\|\|\tilde{s}_{k}-x_{k}\|-\frac{1}{\tau}\|\tilde{s}_{k}-x_{k}\|^{2}\right)$
$\leq\gamma\left(\tau(\|G_{k}^{x}\|+\|\nabla_{x}\mathcal{L}(x_{k},y_{k})\|)\|G_{k}^{x}-\nabla_{x}\mathcal{L}(x_{k},y_{k})\|-\frac{1}{\tau}\|\tilde{s}_{k}-x_{k}\|^{2}\right),$ (28)
where in the first inequality we used Cauchy-Schwartz inequality and the optimality condition of $\tilde{s}_{k}$, and the last inequality follows from nonexpansivity of the projection operator and that $x_{k}=\mathcal{P}_{\mathcal{X}}(x_{k})$. Combining (27) and (5.2) we arrive at
@@
$\gamma\left\langle G_{k}^{x},s_{k}-x_{k}\right\rangle\leq\frac{\gamma}{2}\left(\tau(\|G_{k}^{x}\|+\|\nabla_{x}\mathcal{L}(x_{k},y_{k})\|)\|G_{k}^{x}-\nabla_{x}\mathcal{L}(x_{k},y_{k})\|-\frac{1}{\tau}\|\tilde{s}_{k}-x_{k}\|^{2}-\frac{1}{\tau}\|{s}_{k}-x_{k}\|^{2}\right).$
Since the domains $\mathcal{X}$ and $\mathcal{Y}$ are bounded, and $\mathcal{L}$ is continuously differentiable we conclude that $\|\nabla_{x}\mathcal{L}(x_{k},y_{k})\|$ is bounded. Furthermore, as shown in Lemmas 4.1 and 4.2, the sequences $\{\theta_{k}\}_{k\geq 0}$ and $\{w_{k}\}_{k\geq 0}$ are bounded, which together with continuous differentiability of functions $\Phi$ and $g$ in turn implies the uniform boundedness of $\|G_{k}^{x}\|$. Let us define $M\triangleq\sup_{k\geq 0}\max\{\|G_{k}^{x}\|,\|\nabla_{x}\mathcal{L}(x_{k},y_{k})\|\}<+\infty$. Therefore, plugging the above inequality in (5.2) and rearranging the terms we obtain
@@
$\frac{\gamma}{2\tau}\left\|\tilde{s}_{k}-x_{k}\right\|^{2}$ $\leq f_{\mu}(x_{k})-f_{\mu}(x_{k+1})+\gamma\mathbf{L}_{\mathcal{L}_{2}}\left\|y_{k}-y_{\mu}^{*}(x_{k})\right\|\left\|s_{k}-x_{k}\right\|+2\gamma\tau M\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x}\|$
$\quad+\left(\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma^{2}-\frac{\gamma}{2\tau}\right)\left\|s_{k}-x_{k}\right\|^{2}.$
Summing the above inequality over $k=0$ to $K-1$, dividing both sides by $\gamma K$, and using the second part of Lemma (4.6) we obtain
$\frac{1}{2\tau K}\sum_{k=0}^{K-1}\|\tilde{s}_{k}-x_{k}\|^{2}$ $\leq\frac{f_{\mu}(x_{0})-f_{\mu}(x_{K})}{\gamma K}+\frac{\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}}{K}+\underbrace{\frac{2\tau M}{K}\sum_{k=0}^{K-1}\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x}\|}_{(*)}$
$\quad+\left(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{2}+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma-\frac{1}{2\tau}\right)\frac{1}{K}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}.$ (29)
Next, we establish an upper bound for the term ()(*)( ∗ ). To this end, we begin by obtaining a bound for the consecutive iterates of $y_{k}$ as follows,
$\|y_{k+1}-y_{k}\|$ $\leq\|y_{k}-\sigma\nabla_{y}\mathcal{L}_{\mu}(x_{k},y_{k})-y_{k-1}+\sigma\nabla_{y}\mathcal{L}_{\mu}(x_{k-1},y_{k-1})\|$
$\leq\sigma\|\nabla_{y}\mathcal{L}_{\mu}(x_{k},y_{k})-\nabla_{y}\mathcal{L}_{\mu}(x_{k-1},y_{k})\|+\|y_{k}-\sigma\nabla_{y}\mathcal{L}_{\mu}(x_{k-1},y_{k})-y_{k-1}+\sigma\nabla_{y}\mathcal{L}_{\mu}(x_{k-1},y_{k-1})\|$
$\leq\sigma\mathbf{L}_{\mathcal{L}_{2}}\|x_{k}-x_{k-1}\|+\rho_{D}\|y_{k}-y_{k-1}\|,$
where in the first inequality we used the nonexpansivity of the projection operator, the second inequality is followed by the triangle inequality, and the last inequality is the result of Lipschitz continuity of $\nabla_{y}\mathcal{L}_{\mu}$ and a similar argument as in the proof of Lemma 4.5. Continuing the above recursive relation and using the facts that $\|x_{k+1}-x_{k}\|\leq\gamma\tau M$ and $y_{-1}=y_{0}$, we obtain
$\|y_{k+1}-y_{k}\|\leq\frac{\sigma\gamma\tau}{1-\rho_{D}}\mathbf{L}_{\mathcal{L}_{2}}M.$
Therefore, considering the first part of Lemma 4.8, using the above inequality, and that $\|s_{k}-x_{k}\|\leq\tau M$ for any $k\geq 0$, we obtain
$\sum_{k=0}^{K-1}\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x}\|$ $\leq\Gamma_{4}+\tau M\Gamma_{5}K+C_{\theta x}^{g}\mathbf{C}_{v2}\frac{\sigma\gamma\tau\mathbf{L}_{\mathcal{L}_{2}}}{(1-\rho)(1-\rho_{D})}MK.$
@@
Therefore, using the above inequality within (5.2), the step-size condition $\tau\leq 1/(2\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{2}+(\mathbf{L}_{\mathcal{L}_{1}}+\frac{2\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma)=\mathcal{O}(1/(\kappa_{g}^{3}\gamma))$, and multiplying both sides by $2/\tau$ we obtain
$\frac{1}{\tau^{2}K}\sum_{k=0}^{K-1}\left\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{s}_{k}}-x_{k}\right\|^{2}\leq\frac{2(f(x_{0})-f(x^{*}))+\mu D_{\mathcal{Y}}}{\tau\gamma K}+\frac{2\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+4\tau M\Gamma_{4}}{\tau K}+4\tau\gamma M^{2}\underbrace{\Big{(}\frac{\Gamma_{5}}{\gamma}+\frac{\sigma C_{\theta x}^{g}\mathbf{C}_{v2}\mathbf{L}_{\mathcal{L}_{2}}}{(1-\rho)(1-\rho_{D})}\Big{)}}_{\triangleq B_{4}},$
which completes the first part of the proof by noting that $1-\rho=\Omega(1/\kappa_{g})$, $1-\rho_{D}=\Omega(\mu)$, $\mathbf{L}_{\mathcal{L}_{2}}=\mathcal{O}(\kappa_{g})$, $\mathbf{C}_{v2}=\mathcal{O}(\kappa_{g})$ and $\Gamma_{5}=\mathcal{O}(\gamma\kappa_{g}^{3})$ imply that $B_{4}=\mathcal{O}(\kappa_{g}^{3}/\mu)$.
@@
Next, using the same line of proof as in Theorem 5.1, we can derive a bound for the dual gap function as in (5.1). Therefore, using boundedness of constraint set $\mathcal{X}$ and by noting that $\sum_{i=1}^{k}\rho_{D}^{k-i}\beta^{i}\leq(\frac{\rho_{D}^{k/2}}{1-\beta}+\frac{\beta^{k/2}}{1-\rho_{D}})\beta$ for any $k\geq 2$, we conclude that
@@
Under the premises of Theorem 5.4, let $\mu=\mathcal{O}(\epsilon)$, $\gamma=\mathcal{O}(\epsilon^{3})$ and $\tau=\mathcal{O}(1/\kappa_{g}^{3})$, then for any $K\geq 2\bar{K}=\mathcal{O}(1/\mu\log(1/\mu))$, there exists $t\in\{\bar{K},\ldots,K-1\}$ such that $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$ satisfies $\mathcal{G}_{\mathcal{Z}}(x_{t},y_{t})\leq\epsilon$ within $\mathcal{O}(\frac{\kappa_{g}^{3}}{\epsilon^{5}})$ iterations.
@@
Considering the result of Theorem 5.4 on the primal gap function and multiplying both sides by $K/(K-\bar{K})$ for some $20<\bar{K}\leq K/$ 2. Then, we have that for any $K\geq 2$,
$\frac{1}{\tau^{2}(K-\bar{K})}\sum_{k=\bar{K}}^{K-1}\left\|\tilde{s}_{k}-x_{k}\right\|^{2}$ $\leq\frac{1}{\tau^{2}(K-\bar{K})}\sum_{k=0}^{K-1}\left\|s_{k}-x_{k}\right\|^{2}$
$\leq\frac{2(f(x_{0})-f(x^{*}))+\mu D_{\mathcal{Y}}^{2}}{\tau\gamma(K-\bar{K})}+\frac{2\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+4\tau M\Gamma_{4}}{\tau(K-\bar{K})}+8\tau\gamma M^{2}B_{4}.$
Let $k^{*}=\operatorname*{\arg\!\min}_{k=\bar{K}}^{K-1}\left\|\tilde{s}_{k}-x_{k}\right\|$. Lower-bounding the LHS and taking the square root from both sides of the above inequality, we obtain
$\mathcal{G}_{\mathcal{X}}(x_{k^{*}},y_{k^{*}})=\frac{1}{\tau}\left\|\tilde{s}_{k^{*}}-x_{k^{*}}\right\|$ $\leq\sqrt{\frac{2(f(x_{0})-f(x^{*}))+\mu D_{\mathcal{Y}}^{2}}{\gamma\tau(K-\bar{K})}}+\sqrt{\frac{2\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+4\tau M\Gamma_{4}}{\tau(K-\bar{K})}}+\sqrt{8\tau\gamma M^{2}B_{4}}$
$=\mathcal{O}\left(\frac{1}{\sqrt{\gamma\tau K}}+\sqrt{\frac{\gamma\tau\kappa_{g}^{3}}{\mu}}\right).$